At what temperature (in °C) would one liter of tetrafluoroethylene gas, C2F4, have a mass of 4.192 g, if the gas pressure is 0.8220 atm? Assume that the gas behaves ideally.

________ C
1
Expert's answer
2019-03-18T02:22:59-0400

P(C2F4)×V(C2F4) = n(C2F4) ×R×T(C2F4)

R = 0.082 L×atm×mol-1×K-1

n(C2F4) = m(C2F4) / M(C2F4) = 4.192g / 100g/mol = 0.04192mol.

T(C2F4) = P(C2F4) × V(C2F4) / (n(C2F4) ×R) = 0.8220 atm × 1.0 L / (0.04192 mol × 0.082 L×atm×mol-1×K-1) = 239.1314 K ≈ 239.13 K.

t(C2F4) = T(C2F4) – 273.15 = (239.13 – 273.15)⁰C = –34.02⁰C.